Find the ground-state electron energy by substituting the radial wave function, 𝑅(π‘Ÿ)= 2π‘Žπ‘œ3/2π‘’βˆ’π‘Ÿπ‘Žπ‘œβ„ that corresponds to 𝑛 = 1,𝑙 = 0, into radial equation for hydrogen atom.


1
Expert's answer
2021-01-14T11:03:29-0500

Solution

Given radial wavefunction

"R(r)= (\frac {1}{2a_0})^{\frac {3}{2}} e^{-r/a_0}"

For H- atom enegy is given

"E_n = - \dfrac {m_ee^4}{8 \epsilon ^2_0 h^2 n^2}"

For ground state energy n=1, l=0, m=0

Putting all value then

Ground state energy become

"E_0=-13.6eV"